# Select or Project? Evaluating Lower-dimensional Vectors for LLM Training Data Explanations

### v1

@@ 2 Related Work
Instance-based explanation aims to identify influential training examples that explain a model’s prediction. A dominant framework is leave one out influence (the effect of removing a given training example has on the prediction), which can be approximated with re-training free methods such as influence functions (Koh and Liang, 2017; Grosse et al., 2023). Alternative approaches, such as TracInCP (Pruthi et al., 2020) aim to approximate training data influence during the training process by comparing the test instance’s loss gradient to training gradients at a set of model checkpoints.
@@
These influence estimation methods involve comparisons of model gradients (and calculating inverse Hessian-vector products for influence functions), which remains computationally intractable for LLMs: recent applications therefore restrict computation to a subset of model parameters, such as the multilayer perceptron (MLP) layers (Grosse et al., 2023) or LoRA layers (Kwon et al., 2023). Another common choice is to restrict influence estimation to the model’s final layers (Schioppa et al., 2022; Akyürek et al., 2022; Grosse et al., 2023). However, Yeh et al. (2022) find this to be suboptimal for LLMs due to a cancellation effect in deeper layers, and propose using early layer components instead, for example, the more stable word embedding gradients. Our work extends this idea by systematically comparing the utility of all architectural components to determine which subset provides the most informative signal for retrieval.
@@ 3.4 Random Projection (Project)
Following previous work (e.g., Park et al., 2023; Xia et al., 2024), we create a dense surrogate gradient by projecting the full gradient into a lower-dimensional space using an efficient GPU-optimized implementation from Park et al. (2023). A random projection matrix $\mathbf{R}\in\mathbb{R}^{d\times M}$ (where $d\ll M$) maps the full gradient $\mathbf{v}_{\theta}(s_{i})$ to a smaller vector $\mathbf{p}(s_{i})\in\mathbb{R}^{d}$:

### v2

@@ 2 Related Work
Instance-based explanation aims to identify influential training examples that explain a model’s prediction. A dominant framework is leave one out influence (the effect of removing a given training example has on the prediction), which can be approximated with re-training free methods such as influence functions (Koh and Liang, 2017; Grosse et al., 2023). Alternative approaches, such as TracInCP (Pruthi et al., 2020), aim to approximate training data influence during the training process by comparing the test instance’s loss gradient to training gradients at a set of model checkpoints.
@@
These influence estimation methods involve comparisons of model gradients (and calculating inverse Hessian-vector products for influence functions), which remains computationally intractable for LLMs: recent applications therefore restrict computation to a subset of model parameters, such as the multilayer perceptron (MLP) layers (Grosse et al., 2023) or LoRA layers (Kwon et al., 2023). Another common choice is to restrict influence estimation to the model’s final layers (Schioppa et al., 2022; Akyürek et al., 2022; Grosse et al., 2023). However, Yeh et al. (2022) find this to be suboptimal for LLMs due to a cancellation effect in deeper layers, and propose using early layer components instead, for example, the more stable word embedding gradients. An alternative line of work projects full model gradients to a lower-dimensional space where pairwise distances between examples are preserved (e.g., Park et al., 2023; Xia et al., 2024; Lin et al., 2024). Our work systematically compares the utility of all architectural components and random projection to determine which strategy provides the most informative signal for retrieval.
@@ 3.4 Random Projection (Project)
Following previous work (e.g., Park et al., 2023; Xia et al., 2024; Lin et al., 2024), we create a dense surrogate gradient by projecting the full gradient into a lower-dimensional space using an efficient GPU-optimized implementation from Park et al. (2023). A random projection matrix $\mathbf{R}\in\mathbb{R}^{d\times M}$ (where $d\ll M$) maps the full gradient $\mathbf{v}_{\theta}(s_{i})$ to a smaller vector $\mathbf{p}(s_{i})\in\mathbb{R}^{d}$:
@@ 7 Conclusion
Acknowledgments
This research has been funded by the Vienna Science and Technology Fund (WWTF) [10.47379/VRG19008] "Knowledge-infused Deep Learning for Natural Language Processing". We also thank Lin et al. (2024) for providing the illustration templates of their paper Token-wise Influential Training Data Retrieval for Large Language Models, on which Figure 1 builds.